54007GD:

What is the first step needed to solve 4 over 7 multiplied by x minus 5 equals negative 13 ? (1 point)

Astro:

I am going to also write the equation out to make it easier. \(\dfrac{4}{7}\) \(\times\) (X-5= -13)

Astro:

I made parenethisis just so you didn't get confused between the time sign and the other "X"

Vocaloid:

The way the problem is described I would write it as (4/7)x - 5 = -13 rather than have the parentheses around x - 5. If they wanted to specify the parentheses they would have most likely said “4/ 7 multiplied by the quantity x - 5”. That being said, to solve this, we would first want to isolate (4/7)x. How would we do that? How would we eliminate the -5?

Astro:

@surjithayer wrote:
it can be written both ways. \[\frac{ 4 }{ 7 }(x-5)=13\] or \[\frac{ 4 }{ 7 }x-5=13\]
Remember it's a negative, not a positive.

Vocaloid:

I checked a few other websites w/ the same problem and it seems like they want the 5 outside the parentheses but you are right, the wording of the problem could be less ambiguous

Vocaloid:

we should probably double-check with the OP - is the 5 inside or outside parentheses? @54007GD

mhanifa:

I've seen the same question with answer options: A) Subtract 13 from both sides B) Divide both sides by 7 C) Add 5 to both sides D) Multiply both sides by 4 By analyzing the options we can see A, B and D don't fit either interpretation of the question. A. We would add 13 to both sides B. We would multiply both sides by 7 D. We would divide both sides by 4 We are left with option C. It means the question is: \[\frac{ 4 }{ 7 }x-5=-13\]

maths89898:

Well first, you'd need to multiply as it is the most important

mhanifa:

@maths89898 wrote:
Well first, you'd need to multiply as it is the most important
You would isolate the term with the variable and to do so you need to add 5 to both sides as first step. The next step would be to multiply both sides by 7.

maths89898:

okay